# STATE-NAV: Stability-Aware Traversability Estimation for Bipedal Navigation on Rough Terrain

### v2

@@ IV-A Traversability Estimation via Stability-Aware Command Velocity
where $\operatorname{VaR}(\cdot)$ denotes the Value at Risk [41], calculated from the predicted mean and variance under a Gaussian assumption. This fastest velocity is found approximately by sweeping the command velocity from its maximum candidate $a_{\rm max}=(0.5\;m/s,0.75\;rad/s)$ downward to zero as in Fig. 4, with a resolution $\Delta a=(0.05\;m/s,0.075\;rad/s)$. For terrains where the predicted instability exceeds the threshold even at zero command velocity, we assign minimal values of 0.001 m/s and 0.001 rad/s as the stability-aware command velocities, respectively, to penalize traversal through those regions. Note that $\delta_{\rm limit}$ is a robot-specific hyperparameter that does not need to be adjusted according to environmental conditions, but instead serves to configure the robot’s stability during locomotion. In this paper, we use $\delta_{\rm limit}=3$ and $\alpha=0.97$. Representing traversability as the fastest command velocity under an instability constraint allows the planner to generate navigation-time-optimal paths without weight tuning, while keeping instability within acceptable bounds. This velocity can also be used to constrain the local MPC planner, ensuring stable execution of the high-level navigation plan.
@@ IV-B Traversability-Informed RRT* Global Planner
We propose traversability-informed RRT* (TravRRT*) built upon the works of [42, 23] to generate a global path that minimizes navigation time while aiming to keep instability below a user-defined limit and risk level. During sample generation, the traversability-informed RRT* samples nodes with probability proportional to local traversability, encouraging the search toward more navigable paths. For the cost calculation, the cost $c_{ij}$ of an edge connecting a parent node $n_{i}=(x_{i},y_{i})$ to a child node $n_{j}=(x_{j},y_{j})$ is defined as the expected traversal time along the edge, subject to stability constraints:
@@ IV-C MPC Local Planner
To ensure stable execution of the global path generated by the RRT* planner, we formulate an MPC planner that tracks waypoints while accounting for the reduced-order walking dynamics of Digit. We begin with the Linear Inverted Pendulum (LIP) model [43] to derive discrete dynamics in the local sagittal direction, assuming the fixed step duration $T$. The CoM position change $\Delta x_{q}^{\text{loc}}$ between the $q^{\rm th}$ and ${q+1}^{\rm th}$ walking steps, along with the CoM velocity $v_{q}^{\text{loc}}$, is modeled as a function of the footstep length $u_{q}^{f}$ [10].
@@
where $\hat{\mathbf{x}}_{0}$ is the initial state, and the cost function consists of a terminal cost $m(\mathbf{x}_{N+1})$ which encourages aligning the final heading $\phi_{N+1}$ to the direction to the local waypoint $\phi_{g}$ and reaching the waypoint position $(x_{g},y_{g})$, and the running cost $l(\mathbf{x}_{q},\mathbf{u}_{q})$ which penalizes excessive velocities to promote smooth motion:
@@ V-A Learning Instability with TravFormer
We collect data with the same setting we used for Section III-A. The total simulation time is 8 hours, which corresponds to 72,000 gaitcycles after parsing. With this data, we conduct experiments to evaluate TravFormer’s ability to predict instability with its uncertainty based on elevation maps and command velocities. We design three ablated variants: (i) TravFormer without spatial attention, (ii) TravFormer without deformable attention, and (iii) TravFormer with its transformer decoder replaced by a multi-layer perceptron (MLP) decoder. For fair comparison, each variant selects the checkpoint trained with optimal validation loss in the second phase. We report the root mean squared error (RMSE) and the absolute difference of prediction interval coverage probability ($|\Delta\text{PICP}|$) for all samples, with results shown in Table II.
Table II: RMSE and $|\Delta\text{PICP}|$ results on 3 ablated variants and the full version of TravFormer. $|\Delta\text{PICP}|$ represents the absolute difference between the proportion of ground truth instability values fall within one standard deviation of the predicted range (constructed using predicted instability value $\delta$ and standard deviation $\sigma$, assuming Gaussian distribution) and the ideal coverage of 68.27%. The best values are marked in bold.
@@
The full TravFormer and the two variants that retain the transformer decoder achieve similarly strong performance, outperforming the MLP decoder by 11.8% in RMSE for instability prediction. Furthermore, the full model—which combines spatial attention with deformable self-attention—achieves the best uncertainty estimation among all variants.
Figure 5: Comparison of estimated traversability maps and paths planned by the proposed method and baselines. The starting and goal points are marked by a gray diamond and a yellow star, respectively. (a) Diagonal bird-eye view of the rough terrain environment and the screenshots of the robot traversing along the path by the proposed method. (b) Initialized elevation map of the MuJoCo environment (a). (c), (d), (e) Traversability maps based on (b) and path plans by the QuadFoothold [12], ManualBiped [8], and LearnedInS baselines, respectively. Red and yellow paths are the results of different trade-off weight values $w=3$ and $0.5$. (f) Traversability map based on (b) and the path plan of the proposed method, STATE, with $\delta_{\rm limit}=3$.
Figure 6: Comparison of estimated traversability maps and paths planned by the proposed method and baselines. The starting and goal points are marked by a gray diamond and a yellow star, respectively. (a) Diagonal bird-eye view of the ramp environment and the screenshots of the robot traversing along the path by the proposed method. (b) Initialized elevation map of the MuJoCo environment (a). (c) Traversability map based on (b) and path plans by the LearnedInS baseline. Red and yellow paths are the results of different trade-off weight values $w=3$ and $0.5$. (d) Traversability map based on (b) and the path plan of the proposed method, STATE, with $\delta_{\rm limit}=3$.
@@ V-B1 Implementation
We run the Digit robot in MuJoCo simulation, where point cloud is provided by LiDAR and depth cameras mounted on the robot. The RRT* has an iteration number of 500, and the MPC planner has a horizon length $N$ of 5 walking steps with $w_{g}=1$, $w_{\theta}=5$, and $w_{r}=0.1$. At the beginning of each test, the robot turns around in place at the starting location for 20 seconds to scan the environment and initialize the elevation map. The elevation map and the traversability of the area are kept updated while the robot explores the environment. The whole framework runs on a desktop computer with an Intel Core i9-14900K CPU and NVIDIA GeForce RTX 4090 GPU.
@@ V-B2 Planner Benchmarking
[…]
  • Stability-Aware Traversability Estimation (STATE): The planner utilizes the stability-aware command velocity (5), computed with $\delta_{\rm limit}=3$ and $\alpha=0.97$.
  • Learned Instability (LearnedInS): Traversability estimation is based on the learned risk-sensitive instability output (1), predicted with a maximum linear command velocity input of $(0.5\;m/s,0)$ and $\alpha=0.97$. We take the scaled inverse of the instability to generate a traversability score $t_{k}$, normalized so that flat terrain corresponds to a score of 1, and the higher instability prediction is mapped to a lower score.
  • Manual Traversability Cost Design for Bipeds (ManualBiped) [8]: Traversability estimation uses the manual cost design introduced in [8], which penalizes the slopes and height differences above certain thresholds. Since this design was originally designed to calculate a cost, we apply a scaled inverse transformation in the same manner as in LearnedInS to obtain a traversability score.
  • Quadrupedal Foothold Score (QuadFoothold) [12]: Traversability estimation is based on the foothold quality score proposed in [12], originally developed for indicating valid area that quadrupeds can reliably take its foothold on.
@@ V-B3 Testing Environments
We validate our approach in various simulated environments. We test the algorithms 10 times each in four different 10 m x 10 m MuJoCo environments in Figs. 5, 6, and 7: rough terrain, ramp, undulating terrain, challenging terrain, respectively.
Table III: Navigation results for rough terrain and ramp environments. Success rate is recorded by counting the number of trials that succeeded in reaching the goal. Mean and maximum instability values are taken from the instability recording of each successful trial, and then averaged across the trials. The best values among the planners are marked in bold, and all other values are annotated with the percentage difference relative to the best value.
@@
terrain Mean Instability 2.58 (+4.9%) 2.57 (+4.5%) 2.46 (+0%) 2.79 (+13%) 2.94 (+20%) 3.24 (+32%) 3.12 (+27%)
Max Instability 4.82 (+0%) 7.90 (+64%) 5.37 (+11%) 7.09 (+47%) 8.31 (+72%) 7.50 (+56%) 7.36 (+53%)
@@ V-C1 Benefit of Traversability Learning for Bipedal Locomotion
We first evaluate how well the proposed traversability measures the difficulty in bipedal locomotion in comparison to prior methods. Fig. 5 illustrates the traversability maps and the path plans generated by each planner in the rough terrain scenario. We observe that the QuadFoothold baseline primarily detects risk along the edges of the gray stepping stones, but fails to capture the overall difficulty of the stepped terrain. ManualBiped baseline also underestimates the risk posed by the gray stepped regions, likely due to manually chosen thresholds and hand-crafted rules, which might not be fine-tuned to this type of terrain. Consequently, both planners plan paths that traverse the hazardous areas regardless of the traversability cost weight, which results in lower success rates and higher instability in Table. III. In contrast, our STATE method and the LearnedInS baseline with a high traversability weight leverage our bipedal instability-supervised neural network, assessing stepped areas as more dangerous and successfully avoiding them. These approaches achieve a 100% success rate and lower instability records, demonstrating the advantage of a traversability estimator tailored to bipedal stability.
@@ V-C2 Benefit of Stability-Aware Command Velocity-based Traversability in Global Planning
We then examine the robustness of the proposed global planning strategy by comparing the planning results in different environments. Figs. 5 and 6 depict the path planned by STATE and LearnedInS on rough terrain and ramp environments, respectively. We can observe that STATE plans a safe and efficient path in both environments with a single hyperparameter $\delta_{\rm limit}$. In contrast, the planning results of LearnedInS are highly dependent on environments. On rough terrain, the weight $w=3$ produces a safe and time-efficient path, comparable to that of STATE. However, on ramp, the same weight $w=3$ results in unsafe paths with a lower success rate and higher instability, and even a longer navigation time. This is because the length difference between the shortcut and detour paths is greater in the ramp than that in the rough terrain, necessitating a higher weight to trigger a detour. However, our proposed stability-aware command is environment-agnostic, allowing the planner to choose paths that maintain a relatively consistent instability level.
Figure 7: Path plans of the proposed method and the baselines with the weight $w=3$ in undulating terrain environment. The starting and goal points are marked by a gray diamond and a yellow star, respectively. Green, yellow, blue, and red paths are by the proposed method (STATE), LearnedInS, ManualBiped, and QuadFoothold planners, respectively.
Table IV: Navigation results for undulating terrain environment. Success rate is recorded by counting the number of trials that succeeded in reaching the goal. Mean instability, max instability, and navigation time are calculated for each successful trial and then averaged across the trials.
STATE
LearnedInS
ManualBiped
QuadFoodhold
Success Rate 100% 90% 90% 90%
Mean Instability
2.52
(+17%)
2.30
(+6.5%)
2.16
(+0%)
2.85
(+32%)
Max Instability
4.45
(+4.2%)
4.33
(+1.4%)
4.27
(+0%)
6.73
(+58%)
Navigation Time
39.9 s
(+25%)
46.5 s
(+45%)
57.8 s
(+81%)
32.0 s
(+0%)
@@
Table V: Navigation results for challenging terrain environment. Success rate is recorded by counting the number of trials that succeed in reaching the goal.
@@ V-C3 Benefit of Stability-Aware Command Velocity-based Traversability in MPC Local Planning
Next, we evaluate how stability-aware command velocity improves the execution of global paths in the local MPC planner. Although a global path may be traversable, it often includes regions requiring delicate locomotion, making terrain-aware velocity regulation critical. To evaluate the differences that arise only from local MPC planning behavior, we use the undulating terrain environment with a baseline weight setting of $w=3$, where both the proposed method and baselines generate nearly identical global paths, as shown in Fig. 7. In the final segment of the path, the robot must ascend a moderate slope, demanding careful speed regulation. The navigation results are summarized in Table IV. Note that MPC planners without command constraints, commonly used in prior works, consistently failed and are excluded. In the table, STATE achieved the highest success rate and improved navigation time with only a modest trade-off in stability. This improvement is attributed to a more effective adaptation of motion aggressiveness to terrain conditions: accelerating in favorable terrain and modulating speed in challenging terrain. By contrast, the linear scaling in baselines cannot fully capture the nonlinear relationship between terrain, command, and locomotion stability.
V-C4 Navigation in Challenging Terrain with Fine-grained Roughness
Finally, we present the navigation results in challenging terrain as shown in Fig. 8, demonstrating how the proposed benefits work together to improve navigation performance. The proposed algorithm achieves the highest success rate—though not $100\%$—while maintaining a favorable balance between instability and navigation time shown in Table V. The full simulation results of all four methods are shown in the supplementary video.

### v4

@@ IV-A Traversability as Stability-Aware Command Velocity
where $\operatorname{VaR}(\cdot)$ denotes the Value at Risk [fan2021step], calculated from the predicted mean and variance under a Gaussian assumption. This fastest velocity is found approximately by sweeping the command velocity from its maximum candidate $a_{\rm max}=(0.5\;m/s,0.75\;rad/s)$ downward to zero as in Fig. 3, with a resolution $\Delta a=(0.05\;m/s,0.075\;rad/s)$. For terrains where the predicted instability exceeds the threshold even at zero command velocity, we assign minimal values of 0.001 m/s and 0.001 rad/s as the stability-aware command velocities, respectively, to penalize traversal through those regions. Note that $\delta_{\rm limit}$ is a robot-specific hyperparameter that does not need to be adjusted according to environmental conditions, but instead serves to configure the robot’s stability during locomotion. In this paper, we use $\delta_{\rm limit}=3$ and $\alpha=0.97$. Representing traversability as the fastest command velocity under an instability constraint allows the planner to generate navigation-time-optimal paths without weight tuning, while keeping instability within acceptable bounds. This velocity can also be used to constrain the local MPC planner, ensuring stable execution of the high-level navigation plan.
Remark 1 (Use of risk-sensitive traversability).
We observe that as command velocity increases, the predicted instability grows in a superlinear manner with increasing uncertainty, particularly on steep terrain. By accounting for this behavior, VaR is incorporated to mitigate overestimation of safe velocities and promote stable navigation under uncertainty.
@@ IV-B Traversability-Informed RRT* Global Planner
We propose traversability-informed RRT* (TravRRT*) built upon the works of [karaman2011sampling, chen2024adaptive, gammell2014informed] to generate a global path that minimizes navigation time while aiming to keep instability below a user-defined limit and risk level. During sample generation, the traversability-informed RRT* samples nodes with probability proportional to local traversability, encouraging the search toward more navigable paths. For the cost calculation, the cost $c_{ij}$ of an edge connecting a parent node $n_{i}=(x_{i},y_{i})$ to a child node $n_{j}=(x_{j},y_{j})$ is defined as the expected traversal time along the edge, subject to stability constraints:
@@ IV-C MPC Local Planner
To ensure stable execution of the global path generated by the RRT* planner, we formulate an MPC planner that tracks waypoints while accounting for the reduced-order bipedal walking dynamics. We begin with the LIPM to derive discrete dynamics in the local sagittal direction, assuming the fixed step duration $T$ [gibson2022terrain]. The CoM position change $\Delta x_{q}^{\text{loc}}$ between the $q^{\rm th}$ and ${q+1}^{\rm th}$ walking steps, along with the CoM velocity $v_{q}^{\text{loc}}$, is modeled as a function of the footstep length $u_{q}^{f}$ [shamsah2024terrain].
@@
where $\hat{\mathbf{x}}_{0}$ is the initial state. The cost function includes a terminal cost $m(\mathbf{x}_{N+1})$ which encourages the final heading $\phi_{N+1}$ to align with the direction to the local waypoint $\phi_{g}$ and the position to reach $(x_{g},y_{g})$, and a running cost $l(\mathbf{x}_{q},\mathbf{u}_{q})$ which penalizes excessive velocities to promote smooth motion:
@@ V-A Learning Instability with TravFormer
We collected 72,000 gait cycles data which corresponds to 8 hours simulation time in total for the network training. For evaluation, we tested three ablations: (i) TravFormer without spatial attention, (ii) TravFormer without deformable attention, and (iii) TravFormer with an transformer decoder replaced by an MLP decoder. We report the root mean squared error (RMSE) and the absolute difference of prediction interval coverage probability ($|\Delta\text{PICP}|$) for all samples, with results shown in Table II.
Table II: RMSE and $|\Delta\text{PICP}|$ results on 3 ablated variants and the full version of TravFormer. $|\Delta\text{PICP}|$ represents the absolute difference between the proportion of ground truth instability values fall within one standard deviation of the predicted range (constructed using predicted instability value $\hat{\delta}$ and standard deviation $\sigma$, assuming Gaussian distribution) and the ideal coverage of 68.27%. The best values are marked in bold.
@@
The TravFormer variants with the transformer decoder achieve similarly strong performance, outperforming the MLP decoder by 11.8% in RMSE for instability prediction. Furthermore, the full model—which combines spatial attention with deformable self-attention—achieves the best uncertainty estimation among all variants.
Figure 4: Comparison of estimated traversability maps and paths planned by the proposed method and baselines. The starting and goal points are marked by a gray diamond and a yellow star, respectively. (a) Diagonal bird-eye view of rough terrain environment and the snapshots of the robot traversing along the path by the proposed method. (b) Initialized elevation map of the MuJoCo environment (a). (c), (d), (e) Traversability maps based on (b) and path plans by the QuadFoothold [wellhausen2021rough], ManualBiped [mccrory2023bipedal], and LearnedInS baselines, respectively. Red and yellow paths are the results of different trade-off weight values $w=3$ and $0.5$. (f) Traversability map based on (b) and the path plan of the proposed method, STATE, with $\delta_{\rm limit}=3$.
Figure 5: Comparison of estimated traversability maps and paths planned by the proposed method and baselines. The starting and goal points are marked by a gray diamond and a yellow star, respectively. (a) Diagonal bird-eye view of ramp environment and the snapshots of the robot traversing along the path by the proposed method. (b) Initialized elevation map of the MuJoCo environment (a). (c) Traversability map based on (b) and path plans by the LearnedInS baseline. Red and yellow paths are the results of different trade-off weight values $w=3$ and $0.5$. (d) Traversability map based on (b) and the path plan of the proposed method, STATE, with $\delta_{\rm limit}=3$.
@@ V-B1 Implementation
We run Digit in MuJoCo simulation with point cloud measurements generated by the simulated sensors. The RRT* has an iteration number of 500, and the MPC planner has a horizon length $N$ of 5 walking steps with $w_{g}=1$, $w_{\phi}=5$, and $w_{r}=0.1$. At the beginning of each test, the robot turns around in place at the starting location for 20 seconds to scan the environment and initialize the elevation map. The elevation map and the traversability of the area are kept updated while the robot explores the environment.
@@ V-B2 Planner Benchmarking
To demonstrate the advantages of the proposed method, we design a set of baselines that combine different traversability metrics with various global and local planning strategies. For the cost formulation in global planning, instead of (5), a trade-off between traversability and path length $c_{ij}=\sum_{k\in M}\left(1+w_{t}(1/t_{k})\right)\Delta l_{k}$ is used for baselines, where $t_{k}$ is the traversability score of each approach, and $w_{t}$ is the tunable weight [wermelinger2016navigation]. We test two different weight parameters $w=0.5$ and $3$, to evaluate the effect of the weight on the navigation performance. For local planning, all planners, including the proposed method and the baselines, use the MPC planner introduced in IV-C. Although the baselines do not define traversability in terms of velocity, we extend their MPC implementations to include a comparable velocity constraint as in (14), ensuring a fair comparison. Specifically, each baseline uses its $v^{*},\omega^{*}$ proportional to its respective traversability score, $t_{k}a_{\rm max}$, where $a_{\rm max}=(0.5\;m/s,\;0.75\;rad/s)$ is the maximum velocity used in Section IV-A. In summary, we test the planners as follows:
  • Stability-Aware Traversability Estimation (STATE): The planner utilizes the stability-aware command velocity (5), computed with $\delta_{\rm limit}=3$ and $\alpha=0.97$. $\delta_{\rm limit}=3$ is selected since BSFA instability is roughly 2 for stable forward walking at $0.5$ m/s, so $\delta_{\rm limit}=3$ gives 50% more stability margin for stable walking at 0.5 m/s.
  • Learned Instability (LearnedInS): This method is based on the learned risk-sensitive instability output (1), predicted with a maximum linear command velocity input of (0.5m/s,0)(0.5\;m/s,0) and $\alpha=0.97$. We take the scaled inverse of the instability to generate a traversability score $t_{k}$, normalized so that flat terrain has a score of 1, and the higher instability prediction is mapped to a lower score.
  • Manual Traversability Cost Design for Bipeds (ManualBiped) [mccrory2023bipedal]: This method uses the manual cost design introduced in [mccrory2023bipedal], which penalizes the slopes and height differences above certain thresholds. Since this design was originally designed to calculate a cost, we apply a scaled inverse transformation in the same manner as in LearnedInS to obtain a traversability score.
  • Quadrupedal Foothold Score (QuadFoothold) [wellhausen2021rough]: This method is based on the foothold quality score proposed in [wellhausen2021rough], originally developed for indicating valid area that quadrupeds can reliably take its foothold on.
@@ V-B3 Testing Environments
We validate our approach in various simulated environments. We test the algorithms 10 times each in three 10 m x 10 m MuJoCo environments in Figs. 4-6: rough terrain, ramp, challenging terrain, respectively.
Table III: Navigation results for rough terrain and ramp environments. Success rate is recorded by counting the number of trials that succeeded in reaching the goal. Maximum instability values are taken from the instability recording of each successful trial, and then averaged across the trials. Path smoothness is computed as the mean magnitude of angular acceleration within each successful trial, then averaged across the trials, which increases when the robot changes its heading more often. The best values among the planners are marked in bold, and all other values are annotated with the percentage difference relative to the best value.
@@
terrain Max Instability 4.82 (+0%) 7.90 (+64%) 5.37 (+11%) 7.09 (+47%) 8.31 (+72%) 7.50 (+56%) 7.36 (+53%)
Path Smoothness 0.079 (+0%) 0.119 (+50%) 0.086 (+8.2%) 0.106 (+34%) 0.124 (+56%) 0.114 (+44%) 0.102 (+29%)
@@ V-C1 Benefit of Traversability Learning for Bipedal Locomotion
We first evaluate how well the proposed traversability measures the difficulty in bipedal locomotion in comparison to prior methods. Fig. 4 illustrates the traversability maps and the path plans generated by each planner in rough terrain scenario. We observe that the QuadFoothold baseline primarily detects risk along the edges of the gray stepping stones, but fails to capture the overall difficulty of the region. ManualBiped baseline also underestimates the risk posed by the gray stepped regions, likely due to manually chosen thresholds and hand-crafted rules, which might not be fine-tuned to this type of terrain. Consequently, both planners plan paths that traverse the hazardous areas regardless of the traversability cost weight, which results in lower success rates and higher instability in Table. III. In contrast, our STATE method and the LearnedInS baseline with a high traversability weight leverage our bipedal instability-supervised neural network, assessing stepped areas as more dangerous and successfully avoiding them. These approaches achieve a 100% success rate and lower instability records, demonstrating the advantage of a traversability estimator tailored to bipedal stability.
@@ V-C2 Benefit of Stability-Aware Command Velocity-based Traversability in Global Planning
We then examine the robustness of the proposed global planning strategy by comparing the planning results in different environments. Figs. 4 and 5 depict the path planned by STATE and LearnedInS on rough terrain and ramp environments, respectively. We can observe that STATE plans a safe and efficient path in both environments with a single hyperparameter $\delta_{\rm limit}$. In contrast, the planning results of LearnedInS are highly dependent on environments. On rough terrain, the weight $w=3$ produces a safe and time-efficient path, comparable to that of STATE. However, on ramp, the same weight $w=3$ results in unsafe paths with a lower success rate and higher instability, and even a longer navigation time. This is because the length difference between the shortcut and detour paths is greater in ramp than that in rough terrain, necessitating a higher weight to trigger a detour. However, our proposed stability-aware command is environment-agnostic, allowing the planner to choose paths that maintain a relatively consistent instability level.
Remark 2 (Environment-agnostic hyperparameter tuning of STATE).
For the baseline traversability score definition $t\in[0,1]$ with the cost $c_{ij}=\sum_{k\in M}(1+w(1/t_{k}))^{p}\Delta l_{k}$, the hyperparameters $w$ and $p$ must be re-tuned for each environment, since the score and path length have unrelated scales and units. As a result, in ramp environments, we report that no safe paths are obtained with any $w$ smaller than 100 under $p=1$ (not tested beyond $w=100$), and stable navigation requires the introduction of $p=2$. However, with $p=2$, LearnedInS selects an overly conservative path on rough terrain, resulting in a longer navigation time of 38.3 s. In contrast, tuning $\delta_{\text{limit}}$ in STATE is intuitive and robot-specific, requiring no iterative tuning across environments. These results highlight the main advantage of STATE: defining traversability by stability-aware velocity enables consistent and environment-agnostic path planning without tedious hyperparameter tuning.
@@
Table IV: Navigation results for challenging terrain environment. Evaluation metrics are calculated in the same way as Table. III.
@@ V-C3 Benefit of Stability-Aware Command Velocity-based Traversability in MPC Local Planning
Next, we evaluate how stability-aware command velocity improves the execution of global paths in the local MPC planner. Although a global path may be traversable, it often includes regions requiring delicate locomotion, making terrain-aware velocity regulation critical. To evaluate the differences that arise only from local MPC planning behavior, we use challenging terrain environment with a baseline weight setting of $w=3$, where both the proposed method and LearnedInS generate nearly identical global paths but the paths contain slopes and rough terrains that demand the robot careful speed regulation, as shown in Fig. 6. The navigation results are summarized in Table IV. Note that MPC planners without command constraints, commonly used in prior works, consistently failed and are excluded. The proposed algorithm achieves the highest success rate—though not $100\%$—while maintaining a favorable balance between instability and navigation time shown in Table IV. This improvement is attributed to a more effective adaptation of motion aggressiveness to terrain conditions: accelerating in favorable terrain and modulating speed in unstable terrain. By contrast, the linear scaling in baselines cannot fully capture the nonlinear relationship between terrain, command, and locomotion stability. The full simulation results of all four methods are shown in the supplementary video.
Figure 7: Real-world demonstration of the proposed navigation stack. The left column presents the testing environments with the robot’s executed trajectories overlaid in green. Robot snapshots are extracted from the supplementary video and superimposed for visualization. The right column displays the traversability map and path plan generated by STATE.
V-C4 Real-World Validation
Beyond simulation, we further validate the proposed framework in real-world settings. The robot perceives its environment using a ZED 2i stereo camera mounted on the chest, providing both point cloud perception and localization. All algorithms are run on an external laptop with an Intel Core i7-12700H CPU and an NVIDIA RTX 3060 Mobile GPU. The framework is transferred to hardware with minimal modifications by increasing the MPC horizon length to $7$ and reducing the number of TravRRT* iterations to $400$. The perception and localization modules operate concurrently, with the traversability map and global plan updated every 5 s, and the MPC local planner running at 3 Hz, ensuring real-time performance. We test the system in three outdoor environments, as shown in Fig. 7. With this navigation stack, TravRRT* generates safe navigation paths that avoid untraversable regions identified by TravFormer, while the local MPC planner adaptively slows down when crossing ground obstacles or slopes and accelerates on flat terrain to maintain time efficiency. The robot reaches the goal in $54$, $32$, and $34$ s with maximum BSFA instabilities of $3.4$, $2.0$, and $1.7$, respectively, illustrating the practicality of the framework in real world settings. Full experiments, including active velocity modulation, are provided in the supplementary video.